# Probabilistic Robustness in the Gap Metric

### v1

@@ Abstract
Uncertainties influencing the dynamical systems pose a significant challenge in estimating the achievable performance of a controller aiming to control such uncertain systems. When the uncertainties are of stochastic nature, obtaining hard guarantees for the robustness of a controller aiming to hedge against the uncertainty is not possible. This issue set the platform for the development of probabilistic robust control approaches. In this work, we utilise the gap metric between the known nominal model and the unknown perturbed model of the uncertain system as a tool to gauge the robustness of a controller and formulate the gap as a random variable in the setting with stochastic uncertainties. Main results of this paper includes giving probabilistic bound on the gap exceeding a known threshold followed by bounds on the expected gap value and probabilistic robust stability in terms of the gap metric. Further, we also provide a probabilistic controller performance certification under gap uncertainty and probabilistic guarantee on the achievable $\mathcal{H}_{\infty}$ robustness. Numerical simulations are provided at many places to demonstrate the proposed approach.
I Introduction
@@
The description of perturbed model of the system $\tilde{\Sigma}$ using (1) naturally led several people to conceive the concept of measuring the distance between systems (specifically between Linear Time Invariant (LTI) systems which is of interest to us in this paper). Vidyasagar in [4] proposed the graph metric, followed by gap metric which was proposed in [5]. Since the graph metric was difficult to compute, Georgiou in [6] came up with an elegant formula to compute the gap metric and subsequently the optimal robustness in the gap metric was established in [7]. Meanwhile, Glover in [8] had extended the famous small-gain theorem in [9] to handle perturbation in $\mathcal{L}_{\infty}$ rather than only in $\mathcal{H}_{\infty}$. Vinnicombe leveraged this development and proposed a new metric called the $\nu$-gap in [10] by building upon the gap metric. Authors in [11] proposed a generalised distance measure for LTI systems by taking into account the information about several uncertainty structures. Though in principle, all these metrics can be used for proposing probabilistic robustness, we shall be employing only the gap metric in this paper.
Probabilistic Robust Control (PRC) has been explored before and the interested readers are referred to [12, 13, 14, 15] and the references therein. Specifically for the scenario-based approaches for robust control design, see [16] and the references therein for more details. Having said that, the approach proposed in this paper involves metric between dynamical systems and casting them as random quantities, while the existing PRC approaches define robustness guarantees in terms of the volume of destabilizing perturbations or probability of violation of robust stability and performance conditions. Though our approach presumably aims to solve the same problem, the underlying methodology is certainly different from the existing PRC approaches. In this paper, we shall be using the gap metric for gauging the robustness and use the tools from high dimensional statistics for giving probabilistic guarantees.
I-A Contributions
@@
    We formulate the randomness in the perturbed model $\tilde{\Sigma}$ by resulting it from the stochastic parametric uncertainty for the LTI system and we discuss probabilistic robust stability result in Theorem 7 and we provide probabilistic closed loop guarantees on the residual generalized stability margins for a nominal controller aiming to control the nominal model (See Theorem 8).
@@ I-B Notations
The probability space is defined using a triplet $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega,\mathcal{F}$, and $\mathbb{P}$ denote the sample space, event space, and the probability function, respectively, with $\mathbb{P}:\Omega\rightarrow[0,1]$. A real random vector $x\in\mathbb{R}^{n}$ following a probability density function $\mathbf{f_{x}}$ is denoted by $x\sim\mathbf{f_{x}}$ and its expectation is denoted by $\mathbb{E}[x]$. A zero-mean random vector $x\in\mathbb{R}^{n}$ following a Gaussian distribution with covariance $I_{n}\succ 0$ is denoted by $x\sim\mathcal{N}(0,I_{n})$. Similarly, a random variable $x$ following a Chi-squared distribution with parameter $p>0$ is denoted by $x\sim\chi^{2}_{p}$. A random variable $x\in\mathbb{R}$ is said to be sub-Gaussian with parameter $\sigma>0$ if $\forall\lambda\in\mathbb{R}$, it satisfies $\mathbb{E}[e^{\lambda(X-\mathbb{E}[X])}]\leq e^{\frac{\lambda^{2}\sigma^{2}}{2}}$ or equivalently if $\forall\epsilon>0$, it satisfies $\mathbb{P}(x-\mathbb{E}[x]\geq\epsilon)\leq\exp\left(-\frac{\epsilon^{2}}{2\sigma^{2}}\right)$. The covariance of a random variable $x$ is denoted by $\mathbf{Cov}(x)$.
@@
I-C Preliminaries About Gap Metric
@@
(From [3]) Given $\alpha\in(0,1)$, let the controller $C_{1}$ stabilise the plant model $P_{1}$ and further assume that the pair $(P_{1},C_{1})$ yields a performance measure of $b_{P_{1},C_{1}}>\alpha$, where
$b_{P_{1},C_{1}}$ $:=\left\lVert{\begin{bmatrix}P_{1}\\ I\end{bmatrix}(I-C_{1}P_{1})^{-1}\begin{bmatrix}-C_{1}&I\end{bmatrix}}\right\rVert^{-1}_{\mathcal{H}_{\infty}}.$ (16)
@@
and $\forall P\in\mathcal{B}_{\alpha}(P_{1})$, the performance measure of the controller $C_{1}$ degrades as
@@
Further, from [7] we know that when $\delta_{g}(P,P_{1})<1$,
$\left\lVert{\mathbf{T}_{zw}(P,C_{1})}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{\left\lVert{\mathbf{T}_{zw}(P_{1},C_{1})}\right\rVert_{\mathcal{H}_{\infty}}+\delta_{g}(P,P_{1})}{1-\delta_{g}(P,P_{1})},$ (19a)
$\left\lVert{b_{P,C_{1}}-b_{P_{1},C_{1}}}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{2\delta_{g}(P,P_{1})}{1-\delta_{g}(P,P_{1})}\left\lVert{b_{P_{1},C_{1}}}\right\rVert_{\mathcal{H}_{\infty}}.$ (19b)
The essence of the above proposition is that when the plant model $P_{1}$ is known and when a corresponding stabilising controller $C_{1}$ is also known, it informs us apriori how $C_{1}$ will fair when applied to other plant models in the vicinity of the plant model $P_{1}$ given by plants $P\in\mathcal{B}_{\alpha}(P_{1})$. Particularly, the proposition informs us how the performance measure degrades followed by achievable $\mathcal{H}_{\infty}$ performance degradation in terms of the gap. In the following section, we will formulate the setting where the gap $\delta_{g}(P,P_{1})$ becomes random and that will open a new perspective on doing PRC. Some motivating problems on why the gap $\delta_{g}(P,P_{1})$ may become random and why to develop this theory further are already provided in our earlier work [17].
II Problem Formulation
II-A Uncertain Dynamical System
@@
We will assume that $\theta\sim\mathbf{f_{\theta}}$ where, $\mathbf{f_{\theta}}$ denotes the distribution of the parameter $\theta$. For instance, we can assume that $\mathbf{f_{\theta}}$ is unknown but is believed to be belonging to a moment based ambiguity set $\mathcal{P}^{\theta}$ consistent with mean $\mu_{\theta}\in\mathbb{R}^{p}$ and covariance $\sigma^{2}_{\theta}I_{p}\succ 0$. However, for the ease of exposition, we will assume that $\theta\sim\mathbf{f}_{\theta}=\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$ as this will aid the formulation of the associated gap to become sub-Gaussian which is favourable for obtaining bounds on tail probability (as a Lipschitz function of a Gaussian random variable is sub-Gaussian [20]). Further, we note here that $\tilde{\Sigma}(0)=\bar{\Sigma}$ meaning that the perturbed system with zero uncertainties equals the nominal system (in sense of Figure 1 with $\Delta=0$). This does not imply that $\mu_{\theta}$ is essentially zero as $\mu_{\theta}$ can be non-zero. The only nominal and valid requirement that is needed is that $0\in\mathbf{f}_{\theta}$ (perfectly fine even if the containment happens asymptotically) so that when uncertainties of the perturbed system vanish, it results in the nominal system.
II-B Gap Between Nominal & Perturbed Models
@@
Note that for a fixed $\theta$ value, $\mathrm{Gap}(\theta)$ can be computed using (13). Let $\Theta\subseteq\mathbb{R}^{p}$ denote the set of all possible values of $\theta$ such that, $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$. We now state the main problems of interests that are being addressed in this manuscript.
@@
    bound on the expected value of the $\mathrm{Gap}(\theta)$ denoted by $\mathbb{E}[\mathrm{Gap}(\theta)]$.
@@
Given nominal system model $\bar{\Sigma}$ and the perturbed system model $\tilde{\Sigma}(\theta)$ with $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$ resulting in the associated gap, $\mathrm{Gap}(\theta)$ becoming random, address the following questions:
@@
    find the $\mathbb{E}\left[\left\lVert{\mathbf{T}(\tilde{\Sigma}(\theta))}\right\rVert_{\mathcal{H}_{\infty}}\right]$.
III Solution Methodology
@@
III-A Understanding the Randomness of Graph Operator $\tilde{G}(\theta)$
@@
From assumption 2, we immediately see that the projection operator $\Pi_{\mathcal{G}_{\tilde{\Sigma}(\theta)}}$ also continuously varies with respect to $\theta$. Hence, it’s operator norm would be well-defined. Having mentioned all the assumptions needed for the development of our approach, we will begin our study about giving probabilistic guarantees for robust performance and robust stability in terms of the gap metric.
III-B Probabilistic Guarantee for Coprime Factor Uncertainty
Coprime factor uncertainty can be understood as a combination of multiplicative and inverse multiplicative type uncertainties and the trade off between them is determined by the nominal plant. As a precursor to the gap metric, we will first demonstrate how the randomness in the co-prime factor uncertainty affects the robust stability associated with the nominal controller stabilising the nominal plant. The following theorem (probabilistic extension of Theorem 1 in [7]) gives probabilistic guarantees on the nominal controller stabilising the resulting perturbed plant under random coprime factor uncertainty.
@@ III-B Probabilistic Guarantee for Coprime Factor Uncertainty
III-C Inferring the Lipschitz Constant of $\mathrm{Gap}(\theta)$
@@
III-D Inferring the Expected Value of $\mathrm{Gap}(\theta)$
@@
Under the assumptions of the Lemma 6 and when $\tilde{\Sigma}(0)=\bar{\Sigma}$, we see that
$\mathbb{E}\left[\mathrm{Gap}(\theta)\right]\leq\mathbf{L}_{\mathrm{gap}}\sqrt{\sigma^{2}_{\theta}p+\left\lVert{\mu_{\theta}}\right\rVert^{2}}.$ (39)
@@
Note that $\tilde{\Sigma}(0)=\bar{\Sigma}\iff\mathrm{Gap}(0)=0$. Substituting $\theta^{\prime}=0$ in (36) yields
$\mathbb{E}\left[\mathrm{Gap}(\theta)\right]\leq\mathbf{L}_{\mathrm{gap}}\mathbb{E}\left[\left\lVert{\theta}\right\rVert\right].$ (40)
@@
III-E Probabilistic Robust Stability Guarantees

### v2

@@ Abstract
Uncertainties influencing the dynamical systems pose a significant challenge in estimating the achievable performance of a controller aiming to control such uncertain systems. When the uncertainties are of stochastic nature, obtaining hard guarantees for the robustness of a controller aiming to hedge against the uncertainty is not possible. This issue set the platform for the development of probabilistic robust control approaches. In this work, we utilise the gap metric between the known nominal model and the unknown perturbed model of the uncertain system as a tool to gauge the robustness of a controller and formulate the gap as a random variable in the setting with stochastic uncertainties. The main results of this paper include giving a probabilistic bound on the gap exceeding a known threshold, followed by bounds on the expected gap value and probabilistic robust stability and performance guarantees in terms of the gap metric. We also provide a probabilistic controller performance certification under gap uncertainty and probabilistic guarantee on the achievable $\mathcal{H}_{\infty}$ robustness. Numerical simulations are provided to demonstrate the proposed approach.
{IEEEkeywords}
Probabilistic Robust Control, Gap Metric, Distance Metric
1 Introduction
@@
[…]
The description of perturbed model of the system $\tilde{\Sigma}$ using (1) naturally led several people to conceive the concept of measuring the distance between systems (specifically between Linear Time Invariant (LTI) systems which is of interest to us in this paper). Vidyasagar in [4] proposed the graph metric, followed by gap metric which was proposed in [5]. Since the graph metric was difficult to compute, Georgiou in [6] came up with an elegant formula to compute the gap metric and subsequently the optimal robustness in the gap metric was established in [7]. Meanwhile, Glover in [8] had extended the famous small-gain theorem in [9] to handle perturbation in $\mathcal{L}_{\infty}$ rather than only in $\mathcal{H}_{\infty}$. Vinnicombe leveraged this development and proposed a new metric called the $\nu$-gap in [10] by building upon the gap metric. Authors in [11] proposed a generalised distance measure for LTI systems by taking into account the information about several uncertainty structures. While all these metrics can be used for proposing probabilistic robustness, we shall be employing only the gap metric in this paper. Though gap metric has also been extensively studied for linear time varying (LTV) systems setting too in [12, 13], we shall restrict our study in this manuscript to LTI systems and leave the extension to LTV systems for future work.
Probabilistic Robust Control (PRC) approaches have been explored before and the interested readers are referred to [14, 15, 16, 17] and the references therein. In connection with the scenario-based approaches for robust control design, see [18] and the references therein for more details. Having said that, the approach proposed in this paper involves metric between dynamical systems and casting them as random quantities, while the existing PRC approaches define robustness guarantees in terms of the volume of destabilizing perturbations or equivalently probability of violation of robust stability and performance conditions. Though our approach presumably aims to solve the same problem, the underlying methodology is certainly different from the existing PRC approaches. In this paper, we shall be using the gap metric for gauging the robustness and use the tools from high dimensional statistics for giving probabilistic guarantees. It was already noted by authors in [19] that the gap metric may be not suitable for evaluating the closeness of systems having uncertain poles and zeros on or near the imaginary axis resulting in the difficulty of the stability of the perturbed plant with a stabilizing controller designed for the nominal plant not being guaranteed through the existing stability theorems involving the gap metric. Our developments using probabilistic gap based problem formulation in this manuscript will aim to investigate along with their findings in the sense of probabilistic guarantees.
1.1 Contributions
@@
    We formulate the randomness in the perturbed model $\tilde{\Sigma}$ by resulting it from the stochastic parametric uncertainty for the LTI system and we discuss probabilistic robust stability result in Theorem 7 and we provide probabilistic closed loop deviation (in the $\mathcal{H}_{\infty}$ norm) guarantees due to the action of a nominal controller aiming to control the perturbed model (See Theorem 8).
@@ 1.2 Notations
The probability space is defined using a triplet $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega,\mathcal{F}$, and $\mathbb{P}$ denote the sample space, event space, and the probability function, respectively, with $\mathbb{P}:\Omega\rightarrow[0,1]$. A real random vector $x\in\mathbb{R}^{n}$ following a probability density function $\mathbf{f_{x}}$ is denoted by $x\sim\mathbf{f_{x}}$ and its expectation is denoted by $\mathbb{E}[x]$. A zero-mean random vector $x\in\mathbb{R}^{n}$ following a Gaussian distribution with covariance $I_{n}\succ 0$ is denoted by $x\sim\mathcal{N}(0,I_{n})$. Similarly, a random variable $x$ following a Chi-squared distribution with parameter $p>0$ is denoted by $x\sim\chi^{2}_{p}$. A random variable $x\in\mathbb{R}$ is said to be sub-Gaussian with parameter $\sigma>0$ if $\forall\lambda\in\mathbb{R}$, it satisfies $\mathbb{E}[e^{\lambda(X-\mathbb{E}[X])}]\leq e^{\frac{\lambda^{2}\sigma^{2}}{2}}$ or equivalently if $\forall\epsilon>0$, it satisfies $\mathbb{P}(x-\mathbb{E}[x]\geq\epsilon)\leq\exp\left(-\frac{\epsilon^{2}}{2\sigma^{2}}\right)$. The covariance of a random vector $x\in\mathbb{R}^{n}$ is denoted by $\mathbf{Cov}(x)$.
@@
1.3 Preliminaries About Gap Metric
@@
(From [3]) Given $\alpha\in(0,1)$, let the controller $C_{1}$ stabilise the plant $P_{1}$ and further assume that the pair $(P_{1},C_{1})$ yields a performance measure of $b_{P_{1},C_{1}}>\alpha$, where
$b_{P_{1},C_{1}}$ $:=\left\lVert{Q(P_{1},C_{1})}\right\rVert^{-1}_{\mathcal{H}_{\infty}}\quad\text{and},$ (16a)
$Q(P_{1},C_{1})$ $:=\begin{bmatrix}P_{1}\\ I\end{bmatrix}(I-C_{1}P_{1})^{-1}\begin{bmatrix}-C_{1}&I\end{bmatrix}$ (16b)
@@
and $\forall P\in\mathcal{B}_{\alpha}(P_{1})$, the performance of $C_{1}$ degrades as
@@
Authors in [23] give the following closed loop deviation result
$\left\lVert{Q(P_{1},C_{1})-Q(P,C_{1})}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{\delta_{g}(P,P_{1})}{b_{P,C_{1}}\cdot b_{P_{1},C_{1}}}.$ (19)
Similarly, from [7], we observe that when $\delta_{g}(P,P_{1})<1$,
$\left\lVert{\mathbf{T}_{zw}(P,C_{1})}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{\left\lVert{\mathbf{T}_{zw}(P_{1},C_{1})}\right\rVert_{\mathcal{H}_{\infty}}+\delta_{g}(P,P_{1})}{1-\delta_{g}(P,P_{1})}.$ (20)
The essence of the above proposition is that when the plant model $P_{1}$ is known and when a corresponding stabilising controller $C_{1}$ is also known, it informs us apriori how $C_{1}$ will fair when applied to other plant models in the vicinity of the plant model $P_{1}$ given by plants $P\in\mathcal{B}_{\alpha}(P_{1})$. Particularly, the proposition informs us how the performance measure degrades followed by achievable $\mathcal{H}_{\infty}$ performance degradation in terms of the gap. In the following section, we will formulate the setting where the gap $\delta_{g}(P,P_{1})$ becomes random and that will open a new perspective on doing PRC. Some motivating problems on why the gap $\delta_{g}(P,P_{1})$ may become random and why to develop this theory are available in [20].
2 Problem Formulation
2.1 Uncertain Dynamical System
@@
We will assume that $\theta\sim\mathbf{f_{\theta}}$ where, $\mathbf{f_{\theta}}$ denotes the distribution of the parameter $\theta$. For instance, we can assume that $\mathbf{f_{\theta}}$ is unknown but is believed to be belonging to a moment based ambiguity set $\mathcal{P}^{\theta}$ consistent with mean $\mu_{\theta}\in\mathbb{R}^{p}$ and covariance $\sigma^{2}_{\theta}I_{p}\succ 0$. However, for the ease of exposition, we will assume that $\theta\sim\mathbf{f}_{\theta}=\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$ as this will aid the formulation of the associated gap to become sub-Gaussian which is favourable for obtaining bounds on tail probability (as a Lipschitz function of a Gaussian random variable is sub-Gaussian [24]). Further, we note here that $\tilde{\Sigma}(\theta_{0})=\bar{\Sigma}$ meaning that the uncertainties of the perturbed system vanish at $\theta=\theta_{0}$ and the resulting system equals the nominal system (in sense of Figure 1 with $\Delta=0$). This does not imply that $\mu_{\theta}=\theta_{0}$. The only nominal and valid requirement that is needed is that $\theta_{0}\in\mathbf{f}_{\theta}$ (perfectly fine even if the containment happens asymptotically) so that when uncertainties of the perturbed system vanish, it results in the nominal system. It would be interesting to investigate the randomness in uncertainties by taking into account their structural information as done in [11], but we reserve that research direction for future work. While we don’t take into account the structural information about the uncertainties in this research, we establish guarantees by incorporating further knowledge as to where in the model ambiguity set the system model is more likely to be in the space of (LTI) dynamical systems. We believe that by leveraging this additional knowledge on the space of dynamical systems, we can create new perspective for the field of PRC theory.
2.2 Gap Between Nominal & Perturbed Models
@@
Note that for a fixed $\theta$ value, $\mathrm{Gap}(\theta)$ can be computed using (13). Note that for every realization of $\theta$ say $\bar{\theta}$ from $\mathbf{f}_{\theta}$, we get a deterministic graph subspace $\mathcal{G}_{\tilde{\Sigma}(\bar{\theta})}$ and hence a deterministic projection $\Pi_{\mathcal{G}_{\tilde{\Sigma}(\bar{\theta})}}$. But in general, the randomness in $\tilde{\Sigma}(\theta)$ due to $\theta$ manifests itself as the randomness in the graph subspace $\mathcal{G}_{\tilde{\Sigma}(\theta)}$ and this results in the corresponding projection operator $\Pi_{\mathcal{G}_{\tilde{\Sigma}(\theta)}}$ becoming random as well. Specifically, the projection $\Pi_{\mathcal{G}_{\tilde{\Sigma}(\theta)}}$ is a random operator-valued function of $\theta$. While we certainly want to investigate the above randomness of the projection operator and hence the randomness of the associated $\mathrm{Gap}(\theta)$ using (10), (11), and (12) in detail, we reserve that exciting research direction as a future work. In this manuscript, we will aim to establish the fact that $\mathrm{Gap}(\theta)$ is a sub-Gaussian random variable and leverage the tools from high dimensional statistics to give probabilistic guarantees. Let $\Theta\subseteq\mathbb{R}^{p}$ denote the set of all possible values of $\theta$ such that, $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$. We now state the main problems of interests that are being addressed in this manuscript.
@@
    bound on the $\mathbb{E}[\mathrm{Gap}(\theta)]$.
@@
Given nominal system model $\bar{\Sigma}$ and the perturbed system model $\tilde{\Sigma}(\theta)$ with $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$ resulting in $\mathrm{Gap}(\theta)$ becoming random, address the following questions:
@@
    obtain a bound for $\mathbb{E}\left[\left\lVert{\mathbf{T}(\tilde{\Sigma}(\theta))}\right\rVert_{\mathcal{H}_{\infty}}\right]$.
3 Solution Methodology
@@
3.1 About the Randomness of Graph Operator $\tilde{G}(\theta)$
@@
From assumption 2, we immediately see that the projection operator $\Pi_{\mathcal{G}_{\tilde{\Sigma}(\theta)}}$ also continuously varies with respect to $\theta$. Hence, it’s operator norm would be well-defined. Having mentioned all the assumptions needed, we will begin our study about giving probabilistic guarantees for robust performance and robust stability in terms of the gap metric.
3.2 Probabilistic Guarantee for Coprime Factor Uncertainty
Coprime factor uncertainty can be understood as a combination of multiplicative and inverse multiplicative type uncertainties and the trade off between them is determined by the nominal plant. As a precursor to the gap metric, we will first demonstrate how the randomness in the co-prime factor uncertainty affects the robust stability associated with the nominal controller stabilising the nominal plant. The following theorem (probabilistic extension of Theorem 1 in [7]) gives probabilistic guarantees on the nominal controller stabilising the perturbed plant under random coprime factor uncertainty.
@@ 3.2 Probabilistic Guarantee for Coprime Factor Uncertainty
3.3 Inferring the Lipschitz Constant of $\mathrm{Gap}(\theta)$
@@
3.4 Inferring the Expected Value of $\mathrm{Gap}(\theta)$
@@
Under the assumptions of the Lemma 6 and when $\tilde{\Sigma}(\theta_{0})=\bar{\Sigma}$, we see that
$\mathbb{E}\left[\mathrm{Gap}(\theta)\right]\leq\mathbf{L}_{\mathrm{gap}}\sqrt{\sigma^{2}_{\theta}p+\left\lVert{\mu_{\theta}-\theta_{0}}\right\rVert^{2}}.$ (40)
@@
Note that $\tilde{\Sigma}(\theta_{0})=\bar{\Sigma}\iff\mathrm{Gap}(\theta_{0})=0$. Substituting $\theta^{\prime}=\theta_{0}$ in (37) yields
$\mathbb{E}\left[\mathrm{Gap}(\theta)\right]\leq\mathbf{L}_{\mathrm{gap}}\mathbb{E}\left[\left\lVert{\theta-\theta_{0}}\right\rVert\right].$ (41)
@@
3.5 Probabilistic Robust Stability Guarantees